OpenStudy (anonymous):

Simplified version of x+2/y-3 + 3x-1/y-3.... is the corrrect answer 4x+1/y-3 ??? Medals

jimthompson5910 (jim_thompson5910):

the original expression is this right? \[\Large \frac{x+2}{y+3} + \frac{3x-1}{y+3}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ya thats right

jimthompson5910 (jim_thompson5910):

you have the correct answer assuming you meant to write \[\Large \frac{4x+1}{y+3}\]

jimthompson5910 (jim_thompson5910):

sorry for some reason, I wrote "y+3" when it should be "y-3". My bad

OpenStudy (anonymous):

Thank you both!! And its okay thank you for your help!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

Did i get this one right too? 3/x+1 + 4/x+1.... the answer is 7/x+1 ??

jimthompson5910 (jim_thompson5910):

yes you did

jimthompson5910 (jim_thompson5910):

Nice job

jimthompson5910 (jim_thompson5910):

use parenthesis to indicate you are dividing by all of "x+1" `3/x+1` is NOT the same as `3/(x+1)`

jimthompson5910 (jim_thompson5910):

\[\Large 3/x + 1 = \frac{3}{x} + 1\] \[\Large 3/(x + 1) = \frac{3}{x+1}\]

OpenStudy (anonymous):

Im so sorry I didn't know! I'm new to this so I'm learning

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

Am i correct for this problem also? 5x^2 - x(x-3) = 4x^2- 3x ??

jimthompson5910 (jim_thompson5910):

you distributed incorrectly

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):